# Combining Static and Dynamic Approaches for Mining and Testing Constraints for RESTful API Testing

### v1

@@ 7.5. Ablation Study (RQ5)
In this experiment, we first built a variant that does not contain both observation-confirmation and semantic verifier components. To learn more the static method, the dynamic component was excluded. We modify the prompt in Figure 4 as follows for constraint mining. Instead of providing GPT-4-Turbo with observations derived from another prompt, we feed the description of the parameters and response schema as in the API specification. This prompt replaces Blocks 1- 5 from Figure 3. After providing data on a property, we instruct the LLM to decide if the given property contains a constraint and if there is enough to verify it. This modification merges steps 6 and 7 into a single step. The output is yes or no.
Table 5. Contribution of components in RBCTest (RQ5)
Variant Constraints Mining
@@
RBCTest 435 22 165 95.2 72.5 82.3
RBCTest- 197 127 403 60.8 32.8 42.6
As seen in Table 5, the elimination of observation-confirmation prompting and semantic verifier affects the outcomes, as evidenced by a reduction of 238 correct constraints. Concurrently, there is an increase in the number of false positives. The F1-score of RBCTest- is reduced by half. False positives frequently arise when the model mis-mapping the request parameters with the response properties based on their names. Such mistakes are less common in RBCTest, where the observation prompt is used to enhance the property before it is processed by the confirmation prompt. This result confirms our key contribution of our LLM prompting strategy.
The semantic verifier’s goal is to remove the invalid constraints to improve precision. We currently used a simple verifying mechanism via examples in API specification (Section 2.3). We removed the semantic verifier and ran the static component on two datasets. In the RBCTest dataset with 89 examples, one example invalidated one detected constraint. In the AGORA dataset, with 223 examples over 11 API operations, 6 examples were able to invalidate 6 false-positive constraints. Overall, the verifier accurately confirmed all valid constraints and successfully eliminated 7 incorrect constraints, thus achieving higher precision (89.0% increasing to 91.2%) while maintaining recall (72.5%). These results show that more examples in the API are useful to invalidate more incorrect constraints and confirm the correct ones. Moreover, other types of semantic verifier can be integrated into our framework such as constraint solvers, SMT solvers, domain-specific checkers (valid zip code, phone number format, or valid date checkers), etc.
@@ 8. Threats to Validity
1. Internal Validity. LLM hallucinations might lead to unexpected outcomes. To mitigate this, we incorporated (1) a semantic verifier (Figure 8) and (2) observation-confirmation prompting (Figure 3), which enhance validation through external API specifications and internal consistency checks. Our ablation study confirmed that these methods have a positive impact on performance. We segmented specifications into small sections (Figure 3) to reduce context window, minimizing hallucinations and improving accuracy.
@@
9. Related Work
Recent surveys on API testing (Kim et al., 2022; Golmohammadi et al., 2022; Ehsan et al., 2022; Martin-Lopez et al., 2022; Sharma et al., 2018; Marculescu et al., 2022; Martin-Lopez et al., 2021a) reveal a trend towards automation adoption. AI/ML are used to enhance various aspects of API testing including of generating test cases (Viglianisi et al., 2020; Martin-Lopez et al., 2021b; Atlidakis et al., 2019), realistic test inputs (Alonso et al., 2022), and identify defects early in the development (Liu et al., 2022; Sahin and Akay, 2021; Arcuri, 2020; Arcuri and Galeotti, 2021, 2020; Zhang and Arcuri, 2021a, b; Zhang et al., 2019, 2021). AGORA is a dynamic approach (Alonso et al., 2023).
The advances of LLMs have impacted API testing. (Le et al., 2024) leverages the language capabilities of GPT to fully automate API testing using only an input OpenAPI specification. It builds a dependency graph among API operations, generating test scripts and inputs. Moreover, (Kim et al., 2024) applied GPT to augment specifications, enriching them with explanations of rules and example inputs. Before the era of LLMs, ARTE (Alonso et al., 2022) aimed to generate test inputs for API testing, employing NLP. Morest (Liu et al., 2022) introduced a model-based RESTful API testing method using a dynamically updating RESTful-service Property Graph, showing improvements in code coverage and bug detection.
RESTler (Atlidakis et al., 2019) is a stateful fuzzing tool of REST APIs, which analyzed specifications, inferred dependencies among request types, and dynamically generated tests guided by feedback from service responses. Similarly, RestTestGen (Viglianisi et al., 2020) applied specifications for automatically generating test cases, checking both response status and response data schemas. NLPtoREST (Kim et al., 2023) used the NLP technique Word2Vec (Mikolov et al., 2013) to extract rules from human-readable descriptions in a specification, enhance, and add them back to the specification.
@@ 10. Conclusion
Novelty. This paper presents RBCTest, an approach to mine the constraints on API response bodies and automatically generate test cases to validate them. Our key findings include 1) our LLM-based static method with Observation-Confirmation prompting achieves high precision in constraint mining of 91.2%; and 2) when combining static and dynamic approaches, RBCTest takes advantage of both LLM’s capabilities to comprehend natural language in API specifications, when available, and API execution information to detect constraints on response bodies. With the combination, RBCTest detects more true positive constraints, while reducing a bit precision. RBCTest-generated test cases was able to detect 21 mismatches in real-world APIs, four of which were confirmed by developers in their development forums.
Practical Impact. RBCTest allows teams to test API services of SUTs in both the development and evolution stages. By using API specifications, RBCTest supports applicable use cases where developers need to investigate and test third-party APIs with publicly available specifications before using them for their applications.
Data Availability. Our data and code is publicly available (web, 2025).

### v3

@@ 8.5. Ablation Study (RQ5)
In this experiment, we aim to compare two variants: 1) (RBCTest-): RBCTest without observation-confirmation (OC) prompting and semantic verifier, and 2) RBCTest with both components. To build the former, we modify the prompt for constraint mining in Figure 4 as follows. Instead of providing GPT-4o with observations derived from another prompt, we feed the description of the parameters and response schema as in the API specification. This prompt replaces Blocks 1- 5 from Figure 3. After providing data on a property, we instruct the LLM to decide if the given property contains a constraint and if there is enough to verify it. This modification merges steps 6 and 7 into a single step. The output is yes or no. We ran both variants five times and recorded the mean results in the same process as described in the previous sections.
Table 8. Ablation Study on RBCTest dataset (RQ5)
Variant Constraint Test Generation
@@
RBCTest- 846 279 706 75.2% 54.5% 63.2%
RBCTest 1,137 133 394 89.5% 74.3% 81.2%
As seen in Table 8, the elimination of observation-confirmation prompting and semantic verifier affects the outcomes, as evidenced by a reduction of 312 correct constraints. Concurrently, there is an increase in the number of false positives. The F1-score and recall of RBCTest- decreased by approximately 25% compared to the original RBCTest. False positives frequently arise when the model mis-maps request parameters to response properties based solely on their names. In addition, there are cases where certain fields have names that misleadingly suggest they represent constraints, but based on their descriptions and the predefined evaluation rules, they are actually not constraints. Such mistakes are less common in RBCTest, where the observation prompt is used to enhance the property before it is processed by the confirmation prompt. This result confirms our key contribution of our LLM prompting strategy.
The semantic verifier’s goal is to remove the invalid constraints to improve precision. We currently used a simple verifying mechanism via examples in API specification (Section 4). We removed the semantic verifier and ran the static component on two datasets. In the RBCTest dataset with 89 examples, one example invalidated one detected constraint. In the AGORA dataset, with 223 examples over 11 API operations, 6 examples were able to invalidate 6 false-positive constraints. Overall, the verifier accurately confirmed all valid constraints and successfully eliminated 7 incorrect constraints, thus achieving higher precision (86.4% increasing to 87.5%) while maintaining recall 76.6%. These results show that more examples in the API are useful to invalidate more incorrect constraints and confirm the correct ones. In general, other types of semantic verifier can be integrated into our framework such as constraint solvers, SMT solvers, domain-specific checkers (valid zip code, phone number format, or valid date checkers), etc.
8.6. Token Cost Analysis
For constraint inference, RBCTest does not rely on LLMs to analyze API responses; it instead operates directly on the API specification. This allows each API response field to be processed only once to generate test oracles, which can then be reused across all subsequent API calls. We measured the marginal inference cost of RBCTest with GPT-4o. Due to that strategy, for 19 APIs from both datasets, RBCTest consumed 1,300,479 input tokens (68,445 per API) and 180,527 output tokens (9,501 per API). The total cost for both datasets was \$5.1. 111\$2.5 per 1M input tokens (\$1.25 if cached) \$10 per 1M output tokens.
The test generation phase is more resource-intensive, consuming a total of 1,332,494 input tokens and 473,521 output tokens, with the former accounting for around 73% of the total. The total LLM cost was estimated at \$8.1. The highest token consumption occurred in the GitLab APIs and YouTube_listVideos, with the GitLab Issues API alone using 199,361 input tokens and 70,981 output tokens. Smaller APIs such as Yelp_getBusinesses and OMDB required fewer numbers of tokens.
Overall, the analysis of token usage and monetary cost shows the efficiency of our approach, with a combined cost of \$13.2 across two phases of constraint inference and test script generation for all APIs in the datasets.
@@ 9. Threats to Validity
1. Internal Validity. LLM hallucinations might lead to unexpected outcomes (sallou2024breaking). To mitigate this, we incorporated (1) temperature of zero for GPT-4o, (2) a semantic verifier (Fig. 7) and (3) observation-confirmation prompting (Fig. 3), which enhance validation through external API specifications and internal consistency checks.
Data leakage is expected to be minimal: while specifications and APIs may appear in pre-training data, the specific pairs of specifications and constraints are not. To address non-determinism from input variation, we consistently used the same prompt templates.
@@
10. Discussion on Complementary Nature between Static and Dynamic Approaches
Our goal and empirical results indicate that the static approaches such as RBCTest and the dynamic ones such as AGORA/AGORA+ are complementary to each other.
First, they complement each other in different usage scenarios. The API specifications would be available when 1) API developers provide the OAS specification files for API usages, 2) a tester performs testing based on OAS files, or 3) the front-end developers need to know the specification for back-end integration. If an OAS file is not available, AGORA/AGORA+ could extract the constraints from runtime information, e.g., in the regression testing or other scenarios where the application could be correctly executed. In contrast, if the application is under development and non-executable, RBCTest could rely on the OAS files.
Second, they are also complementary by exploring different information channels. RBCTest detects constraints from API specifications while AGORA/AGORA+ detects invariants through API execution. They provide coverage where the other falls short, either when specifications are incomplete or when execution history is overly specific or unreliable (e.g., during development).
Third, as demonstrated in our results, each approach recovers shared constraints as well as unique ones: RBCTest covers cases missed by runtime test cases or when running is infeasible, while AGORA/AGORA+ handles cases absent from the specification. An example of the complementary result is the one for the Spotify API as mentioned in Section 8.1.4.
Finally, for the mismatch/bug detection application, they are also complementary. RBCTest detects bugs/mismatches between API specification and execution/implementation. AGORA detects bugs/ mismatches across unseen responses in different executions.
@@ 11. Conclusion
Novelty. This work, RBCTest, demonstrates the effectiveness of leveraging LLMs to statically mine logical constraints to be used as oracles for API response bodies directly from API specifications, addressing key limitations of dynamic analysis-based approaches. By introducing the Observation-Confirmation prompting scheme, we reduce hallucinations and significantly improve the precision of mined oracles. RBCTest applies LLMs in combination with API documentation to derive the oracles from the response bodies without relying on runtime execution, enabling broader and more accurate constraint discovery. This novel use of LLMs for static oracle mining represents a shift from traditional behavior-driven testing toward specification-driven validation. RBCTest uncovers 46 real-world inconsistencies between documented specifications and actual APIs.
Practical Impacts. Our results also encourage further exploration into combining static and dynamic techniques—leveraging the strengths of each—to build hybrid systems that can generate richer test oracles, increase coverage, and surface deeper specification-implementation mismatches across software interfaces.
12. Data Availability
Our benchmark and code are publicly available at our website (website).
Acknowledgments
This research is funded by the University of Science, VNU-HCM.